OpenStudy (anonymous):

I'm confused on how to solve this: 54. On a map of downtown, 12th Street is perpendicular to Avenue J. The equation y = —4x + 3 represents 12th Street. What is the equation representing Avenue J if it passes through the point (8, 16)? (from algebra 1 eoc)

OpenStudy (amistre64):

y = 1/4(x-8)+16

OpenStudy (amistre64):

y = (1/4)x +14

OpenStudy (anonymous):

the slope of a perpendicular line is the reciprocal of the slope of the other line and multiplied by a factor of (-1), so the slope of the perpendicular line to to the equation above would be (-1)(-1/4) = 1/4, use this and y-b=m(x-a) and solve giving an answer of y=x/4+14